# Fairness-Regulated Dense Subgraph Discovery

### v1

@@ B.1 Upper bound on the number of binary search iterations for FADSG-I and FADSG-II
Applying binary search to solve problems (1) and (3) requires specifying an initial interval $[L,U]$, where $(L,U)$ denote lower and upper bounds on the optimal value of (1) and (3), respectively, with $L<U$. The number of binary search iterations needed to solve problems (1) and (3) is $\log\frac{U-L}{\Delta},$ with $\Delta$ being the difference of any two distinct values of the objective function of FADSG-I and FADSG-II. It is straight-forward to prove that $\Delta\geq O(1/n^{2})$, similarly to [53, Lemma 1].
B.2 For FADSG-I
@@ B.3 For FADSG-II
Appendix C Tractability of FADSG-II
Our argument is based on a small modification of [47, Lemma 2]. Recall that the numerator of the objective function of problem (3) $p_{2}(\mathcal{S},\lambda)$ is supermodular, but un-normalized with possibly negative values (depending on the value of $\lambda$). In particular, we have $p_{2}(\emptyset,\lambda)=-\lambda|\mathcal{S}_{p}|=-\lambda\cdot n_{p}<0$, since we always select $\lambda>0$. This implies that given a parameter $\beta$, if the quantity $\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)<0$, then $\mathcal{S}\neq\emptyset$. We conclude that any $\mathcal{S}$ for which $\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)<0$ has $p_{2}(\mathcal{S},\lambda)/|\mathcal{S}|>\beta$. Likewise, any $\mathcal{S}$ satisfying $p_{2}(\mathcal{S},\lambda)/|\mathcal{S}|>\beta$ also has $\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)<0$. Overall, we can infer the following equivalence.
$\min_{\mathcal{S}\subseteq\mathcal{V}}\biggl{\{}\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)\biggr{\}}<0\Leftrightarrow\max_{\mathcal{S}\subseteq\mathcal{V}}\biggl{\{}\frac{p_{2}(\mathcal{S},\lambda)}{|\mathcal{S}|}\biggr{\}}>\beta.$ (18)
Next, note that $\beta|\emptyset|-p_{2}(\emptyset,\lambda)=\lambda\cdot n_{p}>0$. Hence, it always holds that $\min_{\mathcal{S}\subseteq\mathcal{V}}\{\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)\}\leq\lambda\cdot n_{p}$. Since the two conditions listed above in (18) are equivalent, their complement is also equivalent. Thus, we obtain
$\min_{\mathcal{S}\subseteq\mathcal{V}}\biggl{\{}\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)\biggr{\}}\in[0,\lambda\cdot n_{p}]\Leftrightarrow\max_{\mathcal{S}\subseteq\mathcal{V}}\biggl{\{}\frac{p_{2}(\mathcal{S},\lambda)}{|\mathcal{S}|}\biggr{\}}\leq\beta.$ (19)
We are now ready to describe the algorithmic strategy for solving (3). Let $\beta_{l},\beta_{u}$ denote lower and upper bounds respectively on the optimal value of (3), and $\beta_{m}:=(\beta_{l}+\beta_{u})/2$ denote the mid-point. Note that the problem
$\min_{\mathcal{S}\subseteq\mathcal{V}}\biggl{\{}\beta_{m}|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)\biggr{\}}$ (20)
[…]
Appendix D Identifying the end of the regularization path
In order to determine the choice of $\lambda$ that corresponds to the end of the regularization path of FADSG-I and II, we can consider analyzing the optimality conditions of problems (1) and (3). Our reasoning behind this step is that for continuous, convex regularization problems, studying the Karush-Kuhn-Tucker (KKT) conditions of the problem (which are necessary and sufficient for optimality) can facilitate the selection of the regularization parameter (see [54]). However, while our formulations are tractable, they are combinatorial in nature, for which there is no analogue of the KKT conditions in general. We note that for the special case where the objective function being maximized is supermodular, the following characterization of the optimality conditions is known.
Fact 1: [48] Consider the problem $\max_{\mathcal{A}\subseteq\mathcal{V}}F(\mathcal{A})$, where $F(\mathcal{A})$ is a supermodular function. Then, a set $\mathcal{A}^{*}$ is optimal if and only if
@@ D.2 FADSG-II
Appendix E Analyzing the Price of Fairness
@@
The above result is pessimistic as it demonstrates that in the worst-case, the price paid in extracting the fair component can be made arbitrarily close to $1$, via a suitable choice of $n$. Additionally, it demonstrates that the lollipop graph is indeed a suitable candidate for analyzing the worst-case trade-off between density and fairness. Next, we present the proof of Proposition 4, in order to understand how the PoF is affected by altering the balance between the majority and minority vertices in the desired soltuion.
E.1 Proof of Proposition 4
@@
Appendix F Super-Greedy++
Super-Greedy++ is a efficient multi-stage greedy algorithm designed to retrieve high-quality solutions for problems referred to as Densest Supermodular Subset (DSS), i.e., maximization of $F(\mathcal{S})/|\mathcal{S}|$, for $F:2^{\mathcal{V}}\rightarrow\mathbb{R}$. This algorithm can extract $(1-\epsilon)$-approximate solutions, for $\epsilon\in(0,1)$ under some conditions in $O(1/\epsilon^{2})$ iterations, as the following result of [17] indicates.
@@ Appendix F Super-Greedy++
Pseudo-code of the algorithm is provided in (Algorithm 1). In order to implement this algorithm for FADSG-I and FADSG-II, one has to derive the expressions of $F(v|\mathcal{S}-v)$, which are provided in Section F.1. For programming simplicity, our implementation at this point does not include heaps. Instead, we adopt a direct approach in which each time we recompute the values of $F(v|\mathcal{S}-v)$ and extract the minimum, while updating vertex degrees after peeling. Therefore, the computational complexity per iteration of our implementation is $O(n^{2})$, resulting in an overall complexity of $O(n^{2}/\epsilon^{2})$.
@@
F.1 Analytical Expressions of $F(v\,|\,\mathcal{S}-v)$.
As defined in the “Introduction” section, $f(v|\mathcal{S}):=f(\mathcal{S}\cup\{v\})-f(\mathcal{S})$. Thus, $F(v\,|\,\mathcal{S}-v)=F(\mathcal{S})-F(\mathcal{S}-v)$, where $\{\mathcal{S}-v\}$ is the set $\mathcal{S}$ without the vertex $v$. Furthermore, with $\deg_{\mathcal{S}}(v)$, we denote the degree of $v$ in the induced set $\mathcal{S}$.
@@
$F(v\,|\,\mathcal{S}-v)=F(\mathcal{S})-F(\mathcal{S}-v)=\left\{\begin{array}[]{ll}\text{deg}_{\mathcal{S}}(v)+\lambda,&\text{if }v\in\mathcal{S}\cap\mathcal{S}_{p}\\ \text{deg}_{\mathcal{S}}(v)-\lambda,&\text{if }v\not\in\mathcal{S}\cap\mathcal{S}_{p}\end{array}\right..$ (54)
Appendix G Dataset details.
@@
Political Books Dataset (PolBooks): The vertices of the network are books on US politics included in the Amazon catalog, with an edge connecting two books if they are frequently co-purchased by the same buyers [^3] https://websites.umich.edu/~mejn/netdata/. Each book is categorized based on its political stance, with possible labels including liberal, neutral, and conservative. In our experiments, we focused solely on the subgraph formed by liberal and conservative books, resulting in $92$ vertices ($43$ associated with a conservative and $49$ with a liberal worldview) connected by a total of $374$ edges, as curated in [21, 22].
Political Blogs Dataset (PolBlogs) [55]: The vertices of the network are weblogs on US politics, with edges representing hyperlinks. Each blog is categorized by its political stance, left or right, with the left ones composing the protected set.
Amazon Products Metadata (Amazon) [56]: The vertices represent Amazon products, and the edges denote frequent co-purchasing product pairs. The attribute associated with each product is their category. In our experiments we used $10$ different such datasets (see Table 3), as assembled in [21, 22]. We name them based on the initials of the categories.
@@
LastFM Asia Social Network (LastFM) [57]: The vertices represent users of LastFM from Asian countries, and the edges are mutual follower connections among them. The attribute associated with each user is their location.
@@
Deezer Europe Social Network (Deezer) [57]: The vertices represent users of Deezer from European countries, and the edges are mutual follower connections among them. The attribute associated with each user is their gender.
GitHub Developers (GitHub) [58]: The vertices represent developers in GitHub who have starred at least $10$ repositories, and the edges are mutual follower relationships between them. The attribute associated with each user is whether they are web or a machine learning developers.
Twitch Users Social Networks (Twitch) [58]: The vertices represent users of Twitch that stream in a certain language, and the edges are mutual friendships between them. The attribute associated with each user is if the user is using mature language.
Twitter Retweet Political Network (Twitter) [59, 60]: The vertices are Twitter users, and the edges represent whether the users have retweeted each other. The attribute associated with each user is their political orientation [61].
Appendix H Details of Experiments
In order to find the appropriate $\lambda$ value for the desired target fairness level that we had for each experiment, we performed bisection on $\lambda$ values. The lower bound was always 00 and the upper was $\lambda_{\max}$. The value of $\lambda_{\max}$ was computed experimentally for each dataset. We specifically chose the value of $\lambda_{\max}$ based on our observation that it led to the densest protected subgraph for FADSG-I and the entire protected subgraph for FADSG-II. Moreover, increasing $\lambda_{\max}$ beyond this point did not alter the desired outcomes. The specific values employed for each dataset and formulation are detailed in Table 6. Furthermore, this bisection was performed with a tolerance on $\lambda$, $\varepsilon=10^{-9}$. Thus, the number of iterations for each bisection experiment was bounded by $\left\lceil\log_{2}\frac{\lambda_{\max}}{\varepsilon}\right\rceil$. Furthermore, in recent work of [53], a stopping criterion for the number of iterations in the Super-Greedy++ algorithm is being introduced, ensuring a certain level of solution’s quality. However, we observed that this criterion tends to be conservative for our purposes. Through experimentation with various datasets, we determined that a value of $T=5$ iterations provided satisfactory results across all our datasets in our implementation of Super-Greedy++. Table 5 shows an example of the running time of each algorithm, on Amazon tmi dataset.
@@ Appendix H Details of Experiments
Table 6: $\lambda_{\max}$ values for each dataset and formulation.
Appendix I Additional Experimental Results
@@ I.1 FADSG-I:
Figures 910 and 11 illustrate the flexibility of our definition of fairness level $\alpha$. They reveal that certain desired fairness levels may be unattainable. As explained in Section 7.1, this limitation arises from the inherent structure of the graph, which does not naturally accommodate certain fairness levels. This can be clearly observed in Figure 10, where the right subfigure exhibits a notable discontinuity just before the curve reaches the value of $0.6$, jumping to a value greater than $0.8$. This is translated in the left subfigure, to an attained fairness level is slightly greater than $0.8$, despite the desired level being $0.6$.
@@
I.2 FADSG-II:
In Figures 13,  14 and 15, we compare FADSG-II with DSG. The top sections show what fraction of all of the protected nodes belongs in the solutions, and the bottom ones depict the PoF for FADSG-II. Again, for some datasets the solution is not exactly the desired one, which is something that can be attributed to the structure of each graph, as it can be seen in Figure 16.

### v2

@@ B.2 Upper bound on the number of binary search iterations for FADSG-I and II
Applying binary search to solve problems (2) and (4) requires specifying an initial interval $[L,U]$, where $(L,U)$ denote lower and upper bounds on the optimal value of (2) and (4), respectively, with $L<U$. The number of binary search iterations needed to solve problems (2) and (4) is $\log\frac{U-L}{\Delta},$ with $\Delta$ being the difference of any two distinct values of the objective function of FADSG-I and II. It is straight-forward to prove that $\Delta\geq O(1/n^{2})$, similarly to Fazzone et al. (2022, Lemma 1).
B.2.1 For FADSG-I
@@ B.2.2 For FADSG-II
Appendix C Identifying the end of the regularization path
In order to determine the choice of $\lambda$ that corresponds to the end of the regularization path of FADSG-I and II, we can consider analyzing the optimality conditions of problems (2) and (4). Our reasoning behind this step is that for continuous, convex regularization problems, studying the Karush-Kuhn-Tucker (KKT) conditions of the problem (which are necessary and sufficient for optimality) can facilitate the selection of the regularization parameter (see Bach et al. (2012)). However, while our formulations are tractable, they are combinatorial in nature, for which there is no analogue of the KKT conditions in general. We note that for the special case where the objective function being maximized is supermodular, the following characterization of the optimality conditions is known.
Fact 1: (Fujishige, 2005) Consider the problem $\max_{\mathcal{A}\subseteq\mathcal{V}}F(\mathcal{A})$, where $F(\mathcal{A})$ is a supermodular function. Then, a set $\mathcal{A}^{*}$ is optimal if and only if
@@ C.2 FADSG-II
Appendix D Analyzing the Price of Fairness
@@
The above result is pessimistic as it demonstrates that in the worst-case, the price paid in extracting the fair component can be made arbitrarily close to $1$, via a suitable choice of $n$. Additionally, it demonstrates that the lollipop graph is indeed a suitable candidate for analyzing the worst-case trade-off between density and fairness. Next, we present the proof of Proposition 7, in order to understand how the PoF is affected by altering the balance between the majority and minority vertices in the desired solution.
D.1 Proof of Proposition 7
@@
D.2 Equivalence of $\delta$ and $\alpha$
As we have already explained in Section 5, the solution of DSG without any fairness considerations is the clique $\mathcal{C}_{\sqrt{n}}$ that contains the unprotected vertices. For this solution, it holds that $r_{2}(\mathcal{S})$ attains its maximum value, which is $r_{2}(\mathcal{S})=\sqrt{n}$. As we decrease the target fairness value $\delta$ of FADSG-II from this upper bound $\delta_{u}=\sqrt{n}$, the solution will gradually include more and more protected vertices from the protected subgraph $\mathcal{P}_{n-\sqrt{n}}$, until we reach the point that we have the whole graph $\mathcal{G}$, for which point it holds that $r_{2}(\mathcal{S})=\frac{1}{\sqrt{n}}$. Up to this point, it always holds that we have a certain fraction (say $11/x<$ 1) of the protected vertices in the solution, i.e., $\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}_{p}|}=\frac{1}{x}<1\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=\frac{n_{p}}{x}=\frac{n-\sqrt{n}}{x}$. This implies that
$r_{1}(\mathcal{S})=\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}|}=\frac{\frac{n-\sqrt{n}}{x}}{\sqrt{n}+\frac{n-\sqrt{n}}{x}}\Leftrightarrow r_{1}(\mathcal{S})=\frac{\sqrt{n}-1}{\sqrt{n}+(x-1)},$ (51)
Moreover, it follows that
$r_{2}(\mathcal{S})$ $=1+\frac{n_{p}}{|\mathcal{S}|}-2r_{1}(\mathcal{S})=1+\frac{n_{p}-2|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}|}=1+\frac{n_{p}-2\frac{1}{x}n_{p}}{|\mathcal{S}|}=1+\frac{\frac{1}{x}(x-2)n_{p}}{|\mathcal{S}|}$ (52)
$=\frac{\sqrt{n}+\frac{x-1}{x}n_{p}}{\sqrt{n}+\frac{1}{x}n_{p}}=\frac{\sqrt{n}+\frac{x-1}{x}(n-\sqrt{n})}{\sqrt{n}+\frac{1}{x}(n-\sqrt{n})}=\frac{1+\frac{x-1}{x}(\sqrt{n}-1)}{1+\frac{1}{x}(\sqrt{n}-1)}=\frac{x+(x-1)\sqrt{n}-x+1}{x+\sqrt{n}-1}$
$\Leftrightarrow r_{2}(\mathcal{S})$ $=\frac{(x-1)\sqrt{n}+1}{\sqrt{n}+(x-1)}$
$\Leftrightarrow x$ $=\frac{(1+r_{2}(\mathcal{S}))(1-\sqrt{n})}{r_{2}(\mathcal{S})-\sqrt{n}},$
which implies that
$r_{1}(\mathcal{S})=\frac{(n+r_{2}(\mathcal{S}))-(r_{2}(\mathcal{S})+1)\sqrt{n}}{n-1}.$ (53)
If we further decrease the target fairness value $\delta$, the solution will contain all of the protected vertices and it is going to start removing the unprotected ones. Hence, it will always hold that $\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}_{p}|}=1\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=n_{p}$. This results in
$r_{2}(\mathcal{S})=1+\frac{n_{p}}{|\mathcal{S}|}-2\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}|}\Leftrightarrow|\mathcal{S}|=\frac{n_{p}-2|\mathcal{S}\cap\mathcal{S}_{p}|}{\delta-1}=\frac{n_{p}}{1-r_{2}(\mathcal{S})}$ (54)
and
$r_{1}(\mathcal{S})=\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}|}=\frac{n_{p}}{\frac{n_{p}}{1-r_{2}(\mathcal{S})}}\Leftrightarrow r_{1}(\mathcal{S})=1-r_{2}(\mathcal{S}).$ (55)
Finally, we get the following mapping of $\delta$ to $\alpha$:
  • if $\delta>\frac{1}{\sqrt{n}}$, with $\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{n_{p}}=\frac{1}{x}<1$, then
    $\alpha=\frac{\sqrt{n}-1}{\sqrt{n}+(x-1)}=\frac{(n+\delta)-(\delta+1)\sqrt{n}}{n-1}$. [corresponds to the 1st branch of Eq. (46)]
  • if $\delta\leq\frac{1}{\sqrt{n}}$, with $\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{n_{p}}=1\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=n_{p}$ , then
    $\alpha=1-\delta$. [corresponds to the 2nd branch of Eq. (46)]
Appendix E Super-Greedy++
Super-Greedy++ is a efficient multi-stage greedy algorithm designed to retrieve high-quality solutions for problems referred to as Densest Supermodular Subset (DSS), i.e., maximization of $F(\mathcal{S})/|\mathcal{S}|$, for $F:2^{\mathcal{V}}\rightarrow\mathbb{R}$. This algorithm can extract $(1-\epsilon)$-approximate solutions, for $\epsilon\in(0,1)$ under some conditions in $O(1/\epsilon^{2})$ iterations, as the following result of Chekuri et al. (2022) indicates.
@@ Appendix E Super-Greedy++
Pseudo-code of the algorithm is provided in (Algorithm 1). In order to implement this algorithm for FADSG-I and II, one has to derive the expressions of $F(v|\mathcal{S}-v)$, which are provided in Section E.1. For programming simplicity, our implementation at this point does not include heaps. Instead, we adopt a direct approach in which each time we recompute the values of $F(v|\mathcal{S}-v)$ and extract the minimum, while updating vertex degrees after peeling. Therefore, the computational complexity per iteration of our implementation is $O(n^{2})$, resulting in an overall complexity of $O(n^{2}/\epsilon^{2})$.
@@
E.1 Analytical Expressions of $F(v\,|\,\mathcal{S}-v)$.
As defined in Section 1, $f(v|\mathcal{S}):=f(\mathcal{S}\cup\{v\})-f(\mathcal{S})$. Thus, $F(v\,|\,\mathcal{S}-v)=F(\mathcal{S})-F(\mathcal{S}-v)$, where $\{\mathcal{S}-v\}$ is the set $\mathcal{S}$ without the vertex $v$. Furthermore, with $\deg_{\mathcal{S}}(v)$, we denote the degree of $v$ in the induced set $\mathcal{S}$.
@@
$F(v\,|\,\mathcal{S}-v)=F(\mathcal{S})-F(\mathcal{S}-v)=\left\{\begin{array}[]{ll}2\text{deg}_{\mathcal{S}}(v)+\lambda,&\text{if }v\in\mathcal{S}\cap\mathcal{S}_{p}\\ 2\text{deg}_{\mathcal{S}}(v)-\lambda,&\text{if }v\not\in\mathcal{S}\cap\mathcal{S}_{p}\end{array}\right..$ (62)
Appendix F Dataset details.
@@
Political Books Dataset (PolBooks): The vertices of the network are books on US politics included in the Amazon catalog, with an edge connecting two books if they are frequently co-purchased by the same buyers [^4] https://websites.umich.edu/~mejn/netdata/. Each book is categorized based on its political stance, with possible labels including liberal, neutral, and conservative. In our experiments, we focused solely on the subgraph formed by liberal and conservative books, resulting in $92$ vertices ($43$ associated with a conservative and $49$ with a liberal worldview) connected by a total of $374$ edges, as curated in Anagnostopoulos et al. (2020; 2024).
Political Blogs Dataset (PolBlogs) (Adamic & Glance, 2005): The vertices of the network are weblogs on US politics, with edges representing hyperlinks. Each blog is categorized by its political stance, left or right, with the left ones composing the protected set.
Amazon Products Metadata (Amazon) (Ni et al., 2019): The vertices represent Amazon products, and the edges denote frequent co-purchasing product pairs. The attribute associated with each product is their category. In our experiments we used $10$ different such datasets (see Table 3), as assembled in Anagnostopoulos et al. (2020; 2024). We name them based on the initials of the categories.
@@
LastFM Asia Social Network (LastFM) (Rozemberczki & Sarkar, 2020): The vertices represent users of LastFM from Asian countries, and the edges are mutual follower connections among them. The attribute associated with each user is their location.
@@
Deezer Europe Social Network (Deezer) (Rozemberczki & Sarkar, 2020): The vertices represent users of Deezer from European countries, and the edges are mutual follower connections among them. The attribute associated with each user is their gender.
GitHub Developers (GitHub) (Rozemberczki et al., 2019): The vertices represent developers in GitHub who have starred at least $10$ repositories, and the edges are mutual follower relationships between them. The attribute associated with each user is whether they are web or a machine learning developers.
Twitch Users Social Networks (Twitch) (Rozemberczki et al., 2019): The vertices represent users of Twitch that stream in a certain language, and the edges are mutual friendships between them. The attribute associated with each user is if the user is using mature language.
Twitter Retweet Political Network (Twitter) (Rossi & Ahmed, 2015; Conover et al., 2011): The vertices are Twitter users, and the edges represent whether the users have retweeted each other. The attribute associated with each user is their political orientation (Tsioutsiouliklis et al., 2021).
Appendix G Details of Experiments
In order to find the appropriate $\lambda$ value for the desired target fairness level that we had for each experiment, we performed bisection on $\lambda$ values. The lower bound was always 00 and the upper was $\lambda_{\max}$. The value of $\lambda_{\max}$ was computed experimentally for each dataset. We specifically chose the value of $\lambda_{\max}$ based on our observation that it led to the densest protected subgraph for FADSG-I and the entire protected subgraph for FADSG-II. Moreover, increasing $\lambda_{\max}$ beyond this point did not alter the desired outcomes. The specific values employed for each dataset and formulation are detailed in Table 6. Furthermore, this bisection was performed with a tolerance on $\lambda$, $\varepsilon=10^{-9}$. Thus, the number of iterations for each bisection experiment was bounded by $\left\lceil\log_{2}\frac{\lambda_{\max}}{\varepsilon}\right\rceil$. Furthermore, in recent work of Fazzone et al. (2022), a stopping criterion for the number of iterations in the Super-Greedy++ algorithm is being introduced, ensuring a certain level of solution’s quality. However, we observed that this criterion tends to be conservative for our purposes. Through experimentation with various datasets, we determined that a value of $T=5$ iterations provided satisfactory results across all our datasets in our implementation of Super-Greedy++. Table 5 shows an example of the running time of each algorithm, on Amazon tmi dataset.
@@ Appendix G Details of Experiments
Table 6: $\lambda_{\max}$ values for each dataset and formulation.
@@
Appendix H Additional Experimental Results
@@ H.1 FADSG-I:
Figures 1213 and 14 illustrate the flexibility of our definition of fairness level $\alpha$. They reveal that certain desired fairness levels may be unattainable. As explained in Section 7.1, this limitation arises from the inherent structure of the graph, which does not naturally accommodate certain fairness levels. This can be clearly observed in Figure 13, where the right subfigure exhibits a notable discontinuity just before the curve reaches the value of $0.6$, jumping to a value greater than $0.8$. This is translated in the left subfigure, to an attained fairness level is slightly greater than $0.8$, despite the desired level being $0.6$.
@@
H.2 FADSG-II:
In Figures 16,  17 and 18, we compare FADSG-II with DSG. The top sections show what fraction of all of the protected nodes belongs in the solutions, and the bottom ones depict the PoF for FADSG-II. Again, for some datasets the solution is not exactly the desired one, which is something that can be attributed to the structure of each graph, as it can be seen in Figure 19.